OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 8x + 2y > 12. A. slope = 4, y-intercept = 6 B. slope = -4, y-intercept = 12 C. slope = negative 1 over 4, y-intercept = 12 D. slope = -4, y-intercept = 6

ganeshie8 (ganeshie8):

it will have same slop & y intercept as the line 8x + 2y = 12 4x + y = 6 y = -4x + 6 slope = -4, y intercept = 6

OpenStudy (anonymous):

thankss :-)

OpenStudy (anonymous):

option 'd' is the correct ans..